# Learning deep illumination-robust features from multispectral filter array images

### v2

@@ IV-C Classification protocol
Training of DNNs. We train using 95% of the training patches, the remaining 5% are used for validation. We use the AdamW weight optimizer for all models. The loss function to be minimized is the cross-entropy loss. Optimization is conducted over 30 epochs with a constant learning rate of $2e$-$4$ and weight decay of $1e$-$5$. The batch size is set to 128.
IV-D Results and discussion
@@
TABLE I: 1-NN classification accuracy (%) on HyTexiLa database with features extracted from either raw or demosaiced test patches. Training/testing are performed according to the protocol described in Sec. IV-C. The best result in each column is shown as bold and second best as italics. Superscript refers to IMEC $4\times 4$, to IMEC $5\times 5$, and to IMEC $2\times 2$.
@@
Input patches Descriptor                                                Feat. size IMEC $4\times 4$ (VIS) IMEC $5\times 5$ (NIR) IMEC $2\times 2$ (VIS-NIR) $64\times 64$ $124\times 124$ $65\times 65$ $125\times 125$ $64\times 64$ $124\times 124$ A Solar A Solar A Solar A Solar A Solar A Solar Raw MSFA RawMixer (ours) 128∗,†,‡ 59.7 83.9 64.8 86.0 80.6 78.4 82.6 80.2 47.4 74.6 56.9 76.5 MSFA-Net [2] 128∗,†,‡ 52.6 81.5 52.2 81.9 79.0 77.5 78.8 75.6 47.9 72.8 52.5 73.7 VGG11raw [12] 512∗,†,‡ 52.0 80.1 54.7 80.8 69.5 68.4 72.4 71.7 52.6 73.0 52.5 71.6 M-LBP [28] 4096/6400/1024 49.4 79.3 49.0 80.7 77.2 73.8 77.7 74.5 23.9 55.8 25.3 58.7 Demosaiced ResNet182 512∗,†,‡ 38.5 93.1 27.7 88.7 50.9 8.4 56.3 33.5 24.4 85.7 29.0 86.2 HSI-Mixer [24] 768∗,†,‡ 17.3 88.6 23.8 87.0 37.4 14.8 39.7 18.4 21.7 75.7 16.7 57.1 ViT [10] 256∗,†,‡ 29.2 87.0 21.8 67.5 31.0 1.7 24.7 6.2 25.3 84.7 17.4 72.6 SpectralFormer [18] 256∗,†,‡ 33.4 86.9 21.7 73.4 38.0 3.0 27.7 8.0 25.3 86.1 21.1 80.1 Vi-Conformer [20] 256∗,†,‡ 31.2 82.9 16.7 73.1 30.7 3.6 25.3 5.7 24.5 85.0 18.0 75.8 Marginal LBP [28] 4096/6400/1024 10.0 41.9 10.5 49.4 26.6 28.5 29.4 33.8 3.0 34.0 2.2 34.1 LAP [9] 256∗,†,‡ 9.4 28.2 9.4 35.5 16.3 18.3 20.9 22.7 7.23 20.9 8.2 25.1 LBP-LCC [8] 512∗,†,‡ 20.0 36.4 22.1 44.8 29.5 22.34 34.9 25.7 6.7 29.1 2.2 32.7
@@
Table I shows the classification results obtained for each feature with 1-nearest neighbor (NN) classifier coupled with the Euclidean distance. Among hand-crafted features, M-LBP outperforms the other LBP-based descriptors since it considers spatio-spectral correlations more effectively within the raw image, while avoiding the demosaicing step that affects the texture representation. Texture features extracted by deep learning-based descriptors from demosaiced patches outperform LBP, LAP, and LBP-LCC features. Their performance is strongly related to both MSFA bands and illuminant. Although WB is performed, the computed features (histograms) have poor discrimination power especially with patches simulated under illuminant A. ResNet18, HSI-Mixer, ViT, SpectralFormer, and Vi-Conformer are outperformed by raw descriptors utilizing IMEC $5\times 5$ patches because they struggle to capture discriminant spatio-spectral interactions from the 25-channel demosaiced patches. However, their performance is significantly enhanced under illuminant A relative to Solar illuminant. This phenomenon can be attributed to the higher RSPD of illuminant A in the NIR bands sampled by IMEC $5\times 5$ filters. ResNet18 and SpectralFormer are ranked first three times and one time among the 12 tested cases, respectively. Our proposed approach is ranked first seven times among the 12 tested cases, followed by MSFA-Net that is ranked second seven times, then VGG11raw, ranked first once and second twice. Unlike RawMixer and MSFANet, VGG11raw struggles to learn consistent spatio-spectral correlations because its first layer is not guided by the MSFA basic pattern. RawMixer exhibits robustness to unseen augmentations and benefits from raw spectral constancy, making its features robust to illumination variation and thus generalizable across different illuminants. Furthermore, learning from raw images by RawMixer provides comparable or superior performance to learning from $B^{2}$-channel demosaiced images at significantly reduced computation costs (see Fig. 6). Its performance is minimally impacted by the MSFA basic pattern size, with similar performance observed under different illuminants (e.g., IMEC $2\times 2$ and IMEC $4\times 4$ under A illuminant, and IMEC $2\times 2$ and IMEC $5\times 5$ under Solar illuminant). This observation shows the significance of both basic pattern size and spectral band arrangement, hinting at the potential extension to the color domain by utilizing different CFAs like the Bayer CFA [26].
@@
(a) (b)
Figure 6: 1-NN classification accuracy vs. computation time of learning feature extraction from the ($N_{l}\approx 27.4\cdot 10^{3}$) $65\times 65$ patches (a), and from the ($N_{l}\approx 8.9\cdot 10^{3}$) $124\times 124$ patches (b). Raw descriptors are highlighted in red, and numbers in parentheses indicate the dimension of each feature. Demosaicing computation time is not considered.
IV-E Ablation study
IV-E1 Effect of raw spectral constancy
To highlight the effect of raw spectral constancy on the classification performance when the illumination conditions change, we reproduce the same experiments described in Section IV without raw spectral constancy. We train RawMixer and MSFA-Net on D65 illuminant raw radiance patches and test on patches associated to illuminants A and Solar with the specific MSFA-preserving augmentations. As we can see from Table II, these experiments show a drastic classification performance decline due to illumination dependency of learned features. This highlights the importance of raw spectral contacy to achieve illumination-robust representation learning.
TABLE II: 1-NN classification accuracy (%) on HyTexiLa raw test patches achieved by RawMixer and MSFANet features following the protocol described in Sec. IV-C. Without raw spectral constancy, learned features become illumination-dependent which drastically alters the texture recognition performance.
@@ A Solar A Solar A Solar A Solar A Solar A Solar
RawMixer 3.9 61.2 3.4 69.8 5.6 5.2 4.8 2.6 2.9 17.8 2.7 20.4
MSFA-Net 3.4 47.4 3.5 61.6 7.6 1.2 8.2 2.1 3.8 12.0 3.2 11.8
TABLE III: 1-NN classification accuracy (%) on HyTexiLa raw test patches achieved by RawMixer and MSFA-Net features following the protocol described in Sec. IV-C. Without MSFA-preserving augmentations, the networks learn with different MSFA patterns which makes learned features inconsistent and less discriminant than those learned from MSFA-preserved patches.
@@
RawMixer 43.8 67.4 9.4 14.4 72.5 71.6 75.0 75.3 43.8 67.4 50.6 72.3
MSFA-Net 39.9 63.8 45.0 67.5 68.2 66.5 73.5 72.3 39.9 63.8 45.0 67.5
IV-E2 Effect of MSFA-preserving augmentation
Our learning from raw images approach implies that the first layer of RawMixer learns spatio-spectral interactions following the basic MSFA pattern that has to be the same for all patches to make learning consistent. To highlight the effect of MSFA pattern preservation during learning, we reproduce the same experiments described in Section IV with augmentations that alter the band arrangement in the basic pattern, using RawMixer and MSFA-Net networks with the same augmentations used in Section IV, however without preserving the MSFA pattern during the augmentation procedure. From Table III, we can clearly see that the performance of features provided by both networks significantly drops when they learn from raw images with altered patterns. Our proposed augmentations ensures texture diversification while preserving the MSFA basic pattern, which makes learned features consistent.
V Conclusion
In this paper, we propose learning illumination-robust features directly from raw images and skip the demosaicing step that alters the spatio-spectral scene representation. We first make raw values less affected by illumination thanks to raw spectral constancy. Then, to diversify raw textures, we propose a strategy for raw image augmentation that preserves the MSFA basic pattern band arrangement. Finally, we propose a model called RawMixer to capture discriminant spatio-spectral interactions in raw images. Extensive experiments on MS image classification with various MSFAs show the relevance of our approach for learning deep illumination-robust texture features from raw images. Future work will focus on improving RawMixer by incorporating raw spectral constancy to its architecture.

### v3

@@ 4.3 Classification protocol
Training of DNNs. We train using 95% of the training patches, the remaining 5% are used for validation. We use the AdamW weight optimizer for all models. The loss function to be minimized is the cross-entropy loss. Optimization is conducted over 30 epochs with a constant learning rate of $2e$-$4$ and weight decay of $1e$-$5$ (training parameters were determined experimentally). The batch size is set to 128.
4.4 Results and discussion
@@
Table 1: 1-NN classification accuracy (%) on HyTexiLa (a) and SpecTex (b) datasets with features extracted from either raw or demosaiced test patches. Training/testing are performed according to the protocol described in Sec. 4.3. The best result in each column is shown as bold and second best as italics. Superscript refers to IMEC $4\times 4$, to IMEC $5\times 5$, and to IMEC $2\times 2$. Note that IMEC $5\times 5$ is not considered for SpecTex dataset since it’s not defined in the NIR domain.
@@
Input patches Descriptor                                                Feat. size IMEC $4\times 4$ (VIS) IMEC $5\times 5$ (NIR) IMEC $2\times 2$ (VIS-NIR) $64\times 64$ $124\times 124$ $65\times 65$ $125\times 125$ $64\times 64$ $124\times 124$ A Solar A Solar A Solar A Solar A Solar A Solar Raw MSFA RawMixer (ours) 128∗,†,‡ 59.7 83.9 64.8 86.0 80.6 78.4 82.6 80.2 47.4 74.6 56.9 76.5 MSFA-Net [2] 128∗,†,‡ 52.6 81.5 52.2 81.9 79.0 77.5 78.8 75.6 47.9 72.8 52.5 73.7 VGG11raw [13] 512∗,†,‡ 52.0 80.1 54.7 80.8 69.5 68.4 72.4 71.7 52.6 73.0 52.5 71.6 M-LBP [29] 4096/6400/1024 49.4 79.3 49.0 80.7 77.2 73.8 77.7 74.5 23.9 55.8 25.3 58.7 Demosaiced ResNet182 512∗,†,‡ 38.5 93.1 27.7 88.7 50.9 8.4 56.3 33.5 24.4 85.7 29.0 86.2 HSI-Mixer [25] 768∗,†,‡ 17.3 88.6 23.8 87.0 37.4 14.8 39.7 18.4 21.7 75.7 16.7 57.1 ViT [11] 256∗,†,‡ 29.2 87.0 21.8 67.5 31.0 1.7 24.7 6.2 25.3 84.7 17.4 72.6 SpectralFormer [19] 256∗,†,‡ 33.4 86.9 21.7 73.4 38.0 3.0 27.7 8.0 25.3 86.1 21.1 80.1 Vi-Conformer [21] 256∗,†,‡ 31.2 82.9 16.7 73.1 30.7 3.6 25.3 5.7 24.5 85.0 18.0 75.8 Marginal LBP [29] 4096/6400/1024 10.0 41.9 10.5 49.4 26.6 28.5 29.4 33.8 3.0 34.0 2.2 34.1 LAP [10] 256∗,†,‡ 9.4 28.2 9.4 35.5 16.3 18.3 20.9 22.7 7.23 20.9 8.2 25.1 LBP-LCC [9] 512∗,†,‡ 20.0 36.4 22.1 44.8 29.5 22.3 34.9 25.7 6.7 29.1 2.2 32.7
(a) HyTexila dataset
Input patches Descriptor                                                Feat. size IMEC $4\times 4$ (VIS) IMEC $2\times 2$ (VIS-NIR) $64\times 64$ $124\times 124$ $64\times 64$ $124\times 124$ A Solar A Solar A Solar A Solar Raw MSFA RawMixer (ours) 128∗,‡ 83.1 91.4 82.6 91.5 77.9 84.6 79.5 85.2 MSFA-Net [2] 128∗,‡ 79.2 89.0 77.8 88.6 73.5 83.5 77.2 84.5 VGG11raw [13] 512∗,‡ 74.9 82.5 64.0 70.7 72.1 79.6 67.4 73.4 M-LBP [29] 4096/1024 73.2 89.3 73.4 90.7 42.3 69.5 42.2 69.5 Demosaiced ResNet182 512∗,‡ 29.8 92.1 21.5 85.1 63.6 59.1 28.8 63.6 HSI-Mixer [25] 768∗,‡ 25.2 89.1 26.9 84.8 46.3 58.0 30.6 46.3 ViT [11] 256∗,‡ 21.1 83.9 15.7 62.7 50.5 59.7 13.6 49.1 SpectralFormer [19] 256∗,‡ 32.0 92.7 23.3 72.9 49.2 63.1 15.7 49.2 Vi-Conformer [21] 256∗,‡ 22.2 82.3 19.2 55.1 49.1 60.5 13.2 50.5 Marginal LBP [29] 4096/1024 50.1 72.0 41.7 67.2 16.0 52.7 16.2 47.3 LAP [10] 256∗,‡ 31.8 46.3 34.6 60.6 26.7 38.9 29.9 39.4 LBP-LCC [9] 512∗,‡ 39.7 60.3 47.7 70.5 40.2 57.9 38.2 52.1
(b) SpecTex dataset
@@
Table 1(b) shows the classification results obtained for each feature with 1-nearest neighbor (NN) classifier coupled with the Euclidean distance. On HyTexila dataset (see Tab. 1(b)(a)), M-LBP outperforms the other LBP-based descriptors since it considers spatio-spectral correlations more effectively within the raw image, while avoiding the demosaicing step that affects the texture representation. Texture features extracted by deep learning-based descriptors from demosaiced patches outperform LBP, LAP, and LBP-LCC features. Their performance is strongly related to both MSFA bands and illuminant. Although WB is performed, the computed features (histograms) have poor discrimination power especially with patches simulated under illuminant A. ResNet18, HSI-Mixer, ViT, SpectralFormer, and Vi-Conformer are outperformed by raw descriptors utilizing IMEC $5\times 5$ patches because they struggle to capture discriminant spatio-spectral interactions from the 25-channel demosaiced patches. However, their performance is significantly enhanced under illuminant A relative to Solar illuminant. This phenomenon can be attributed to the higher RSPD of illuminant A in the NIR bands sampled by IMEC $5\times 5$ filters. ResNet18 and SpectralFormer are ranked first three times and one time among the 12 tested cases, respectively. They perform better under Solar illuminant in the VIS and VIS-NIR domains, likely because textures are more prominently highlighted in these conditions. However, their performance declines compared to RawMixer and MSFA-Net, possibly due to the artifacts introduced by the demosaicing procedure, which can disrupt texture representation. Additionally, capturing spatio-spectral correlations from demosaiced images is challenging, even with advanced learning paradigms. Replacing the 2D convolutional kernels in some layers with 3D kernels could improve their performance, as 3D kernels capture both spatial and spectral information simultaneously. Our proposed approach is ranked first seven times among the 12 tested cases, followed by MSFA-Net that is ranked second seven times, then VGG11raw, ranked first once and second twice. Unlike RawMixer and MSFANet, VGG11raw struggles to learn consistent spatio-spectral correlations because its first layer is not guided by the MSFA basic pattern. On the SpecTex dataset (see Tab. 1(b)(b)), RawMixer ranks first in seven out of the eight tested cases, followed by MSFA-Net, which ranks second in six cases. While VGG11raw performs better than M-LBP with IMEC $2\times 2$, M-LBP features outperform those of VGG11raw with IMEC $4\times 4$, likely because M-LBP captures more spatio-spectral channel correlations with this MSFA. Additionally, M-LBP surpasses the other tested LBP-based descriptors. Similar to the HyTexila results, ResNet18, SpectralFormer, and HSI-Mixer deliver strong performances in the VIS domain under Solar illuminant, likely because textures are better highlighted with IMEC $4\times 4$ under Solar conditions. RawMixer exhibits robustness to unseen augmentations and benefits from raw spectral constancy, making its features robust to illumination variation and thus generalizable across different illuminants and textures. Furthermore, learning from raw images by RawMixer provides comparable or superior performance to learning from $B^{2}$-channel demosaiced images at significantly reduced computation costs (see Fig. 7). Its performance is minimally impacted by the MSFA basic pattern size, with similar performance observed under different illuminants (e.g., IMEC $2\times 2$ and IMEC $4\times 4$ under A illuminant, and IMEC $2\times 2$ and IMEC $5\times 5$ under Solar illuminant), hinting at the potential extension to the color domain by utilizing different CFAs like the Bayer CFA [27].
@@
(a) IMEC $5\times 5\mid 65\times 65\mid N_{l}\approx 27.4\cdot 10^{3}$ (b) IMEC $4\times 4\mid 124\times 124\mid N_{l}\approx 38.4\cdot 10^{2}$
Figure 7: 1-NN classification accuracy vs. computation time of feature extraction from the learning patches associated to HyTexila (a) and SpecTex (b) datasets. Raw descriptors are highlighted in red, and numbers in parentheses indicate the dimension of each feature. Demosaicing computation time (e.g., $\approx$41s/image for HyTexila images with IMEC $5\times 5$ MSFA) is not considered.
4.5 Ablation study
4.5.1 Effect of raw spectral constancy
To highlight the effect of raw spectral constancy on the classification performance when the illumination conditions change, we reproduce the same experiments described in Section 4 using RawMixer but without raw spectral constancy. As we can see from Table 2, these experiments show a drastic classification performance decline due to illumination dependency of learned features. This highlights the importance of raw spectral constancy to achieve illumination-robust representation learning.
Table 2: 1-NN classification accuracy (%) on HyTexiLa (first row) and SpecTex (second row) raw test patches achieved by RawMixer features following the protocol described in Sec. 4.3. Numbers in parentheses are the results (without ablation) recalled from Table 1(b). Without raw spectral constancy, learned features become illumination-dependent which drastically alters the texture recognition performance.
@@ A Solar A Solar A Solar A Solar A Solar A Solar
HyTexila 3.9 (59.7) 61.2 (83.9) 3.4 (64.8) 69.8 (86.0) 5.6 (80.6) 5.2 (78.4) 4.8 (82.6) 2.6 (80.2) 2.9 (47.4) 17.8 (74.6) 2.7 (56.9) 20.4 (76.5)
SpecTex 8.0 (83.1) 73.3 (91.4) 6.0 (82.6) 74.2 (91.5) - - 2.3 (77.9) 24.5 (84.6) 4.6 (79.5) 27.8 (85.2)
4.5.2 Effect of MSFA-preserving augmentation
Our learning from raw images approach implies that the first layer of RawMixer learns spatio-spectral interactions following the basic MSFA pattern that has to be the same for all patches to make learning spatio-spectral interactions consistent. To highlight the effect of MSFA pattern preservation during learning, we repeated the experiments from Section 4 with RawMixer using the same augmentations as before but without preserving the MSFA pattern during augmentation. Table 3 clearly shows that the performance of features provided by the network significantly drops when it learns from raw images with altered patterns, especially on HyTexila dataset. Our proposed augmentations ensures texture diversification while preserving the MSFA basic pattern, which makes learned features consistent.
Table 3: 1-NN classification accuracy (%) on HyTexiLa (first row) and SpecTex (second row) raw test patches achieved by RawMixer features following the protocol described in Sec. 4.3. Numbers in parentheses are the results (without ablation) recalled from Table 1(b). In this case, the network learns varying MSFA patterns, leading to inconsistencies and less discriminant features, especially with HyTexiLa patches.
@@
HyTexila 43.8 (59.7) 67.4 (83.9) 9.4 (64.8) 14.4 (86.0) 72.5 (80.6) 71.6 (78.4) 75.0 (82.6) 75.3 (80.2) 43.8 (47.4) 67.4 (74.6) 50.6 (56.9) 72.3 (76.5)
SpecTex 81.9 (83.1) 89.3 (91.4) 82.4 (82.6) 89.4 (91.5) - - 77.0 (77.9) 84.2 (84.6) 78.6 (79.5) 84.8 (85.2)
5 Conclusion
In this paper, we propose learning illumination-robust features directly from raw images, bypassing the demosaicing step that alters the spatio-spectral scene representation. We introduce raw spectral constancy to minimize illumination effects and a raw image augmentation strategy that preserves the MSFA band arrangement. Finally, we propose a model called RawMixer to capture discriminant spatio-spectral interactions in raw images. Extensive experiments on MS image classification show the relevance of our approach for learning deep, illumination-robust texture features. Future work will focus on enhancing RawMixer by integrating raw spectral constancy into its architecture and developing a specialized encoder for MS image segmentation. Potential applications include outdoor crop/weed recognition.